OpenStudy (anonymous):

If g(x)=2x+1 and h(x)=4x^2+4x+7, find a function f such that f(g(x)=h(x). The answer is f(x)=x^2+6

OpenStudy (anonymous):

You need to rewrite H(x) so that you can see g(x) in it h(x) = (2x + 1) ^2 + 6 g(x) = 2x + 1 so therefore, F(x) = x^2 + 6

OpenStudy (anonymous):

Can you show the step?